Draw the following triangle: A (-3, 2), B (9, 2), and C (3, -6) and use it to answer the following question.
perimeter =

Answer by Jk22(389)


perimeter = topline+2*sideline(pythagoras)= 12+2*sqrt(6*6+8*8)=12+2*10=32